OpenStudy (anonymous):

I know you can find this problem online anywhere, but I am just practicing it.

OpenStudy (anonymous):

\[\large \int\limits_{ }^{ } \frac{e^x}{x}~dx\]we integrate by parts, differentiating the x on the bottom, and integrating e^x.\[\large \int\limits_{ }^{ } \frac{e^x}{x}~dx= \frac{e^x}{x}-\large \int\limits_{ }^{ } \frac{e^x}{-x^2}~dx\]\[\large \int\limits_{ }^{ } \frac{e^x}{x}~dx= \frac{e^x}{x}+\large \int\limits_{ }^{ } \frac{e^x}{x^2}~dx\]then same integration by parts,\[\large \int\limits_{ }^{ } \frac{e^x}{x}~dx= \frac{e^x}{x}+ \frac{e^x}{x^2}~dx-\large \int\limits_{}^{ }\frac{-2e^x}{x^3}dx \]\[\large \int\limits_{ }^{ } \frac{e^x}{x}~dx= \frac{e^x}{x}+ \frac{e^x}{x^2}~dx+\large \int\limits_{}^{ }\frac{2e^x}{x^3}dx \]and on,\[\large \int\limits_{ }^{ } \frac{e^x}{x}~dx= \frac{e^x}{x}+ \frac{e^x}{x^2}+\frac{2e^x}{x^3} +\large \int\limits_{}^{ }\frac{6e^x}{x^4}dx \]\[\large \int\limits_{ }^{ } \frac{e^x}{x}~dx= \frac{e^x}{x}+ \frac{e^x}{x^2}+\frac{2e^x}{x^3} +\frac{6e^x}{x^4}+\large \int\limits_{}^{ } \frac{24e^x}{x^5}dx \]\[ \int\limits_{ }^{ } \frac{e^x}{x}~dx= \frac{e^x}{x}+ \frac{e^x}{x^2}+\frac{2e^x}{x^3} +\frac{6e^x}{x^4}+ \frac{24e^x}{x^5} +\large \int\limits_{}^{ } \frac{120e^x}{x^6}dx\]\[ \int\limits_{ }^{ } \frac{e^x}{x}~dx= \color{red}{\frac{e^x}{x}+ \frac{e^x}{x^2}+\frac{2e^x}{x^3} +\frac{6e^x}{x^4}+ \frac{24e^x}{x^5} + \frac{120e^x}{x^6}...}\]

OpenStudy (anonymous):

I did the integration part, but now I want to think of writing it as a power series. \[ \int\limits_{ }^{ } \frac{e^x}{x}~dx= \color{red}{\left(\begin{matrix} \\ \frac{e^x}{x}+ \frac{e^x}{x^2}+\frac{2e^x}{x^3} +\frac{6e^x}{x^4}+ \frac{24e^x}{x^5} + \frac{120e^x}{x^6}... \end{matrix}\right)}\]\[ \int\limits_{ }^{ } \frac{e^x}{x}~dx=~ \frac{e^x}{x}\color{red}{\left(\begin{matrix} \\ 1+ \frac{e^x}{x}+\frac{2}{x^3} +\frac{6}{x^2}+ \frac{24}{x^4} + \frac{120}{x^5}... \end{matrix}\right)}\]

OpenStudy (anonymous):

if I am doing anything wrong, then please correct me....

OpenStudy (anonymous):

I am off at the last line

OpenStudy (anonymous):

\[ \int\limits_{ }^{ } \frac{e^x}{x}~dx=~ \frac{e^x}{x}\color{red}{\left(\begin{matrix} \\ 1+ \frac{1}{x}+\frac{2}{x^3} +\frac{6}{x^2}+ \frac{24}{x^4} + \frac{120}{x^5}... \end{matrix}\right)}\]corrected.

OpenStudy (anonymous):

I will refersh.

OpenStudy (anonymous):

now, I need to do think of a sigma for this whole red thing. \[\large \sum_{n=0}^{\infty}~\color{red}{ (n+1)x^{-n}}\]

OpenStudy (anonymous):

my powers in the red, where I have integral of e^x/x dx= ... are disordered, it should be x^2 on the bottom and then x^3. but I get the idea, do I ?

OpenStudy (anonymous):

\[ \int\limits_{ }^{ } \frac{e^x}{x}~dx=~ \frac{e^x}{x} \times \left(\begin{matrix} \sum_{n=0}^{\infty}~n(n+1)(x^{-n})\\ \end{matrix}\right)\]

OpenStudy (anonymous):

error was in the sigma as well. I needed 1 * (1+0) * x^0 2 * (1+1) * x^-1 3 * (1+2) * x^-2

OpenStudy (anonymous):

I am done....

OpenStudy (anonymous):

if any corrections, besides ones that I have caught myself on.... if not, then over I guess.

OpenStudy (xapproachesinfinity):

the summation you did should be looked at again doesn't seem to me all correct

OpenStudy (xapproachesinfinity):

i believe you missed signs as well in the integration by part! let me see that

OpenStudy (anonymous):

are you sure? can you tell me where I made a mistake please? it seems as though, that we get: (1)*(1+0)*x^(-0) --> 1 . when n=0 (1)*(1+1)*x^(-1) --> 2*x^-2 . when n=1 (2)*(1+2)*x^(-2) --> 6*x^-3.. when n=2 (3)*(1+3)*x^(-3) --> 24*x^-4. when n=3

OpenStudy (xapproachesinfinity):

for n=1 spot that you have 2/x not 2/x^2

OpenStudy (xapproachesinfinity):

for the sign i take back seems fine^_^

OpenStudy (anonymous):

you mean that for n=1, I get 2/x instead of 1/x maybe?

OpenStudy (anonymous):

still a mistake, but it would be better to know where it is

OpenStudy (xapproachesinfinity):

yes

OpenStudy (anonymous):

but the integration part is alright?

OpenStudy (anonymous):

I'll be typing for the sigma purposes for now...

OpenStudy (xapproachesinfinity):

yes seems good to me! you just need to rewrite the summation correctly

OpenStudy (anonymous):

\[\frac{e^x}{x}\left(\begin{matrix} \\ \end{matrix}1+ \frac{1}{x}+\frac{2}{x^2}+\frac{6}{x^3}+\frac{24}{x^4}+\frac{120}{x^5} \right)\]now once I seee it I will think

OpenStudy (anonymous):

yeah, even the n(n+1) part was wrong

OpenStudy (anonymous):

I see the pattern so I should be able to figure out the sigma

OpenStudy (xapproachesinfinity):

yes! i believe so :)

OpenStudy (anonymous):

\[\large \sum_{n=0}^{\infty } \frac{\rm blank~~for~~now}{x^n}\]

OpenStudy (xapproachesinfinity):

yes! the top should be decrease in power compared to the bottom so even if you see the pattern you need to see this part as well

OpenStudy (anonymous):

ohh, it involves factorial!!! 0!=1 1!=1 2!=2 3!=6 4!=24 5!=120 and on,.....!! omg

OpenStudy (xapproachesinfinity):

yes there is factorial indeeed

OpenStudy (xapproachesinfinity):

i guess your summation becomes obvious now^_^

OpenStudy (anonymous):

so, lets do this. \[\frac{e^x}{x}\left(\begin{matrix} \\ \end{matrix} \sum_{n=-1}^{\infty} \frac{(n+1)!}{x^{n+1}} \right)\]

OpenStudy (anonymous):

am I allowed to start from n=-1?

OpenStudy (xapproachesinfinity):

why would you do that?

OpenStudy (anonymous):

to have 0! for first term, and 1! for the second term.

OpenStudy (xapproachesinfinity):

from zero is fine

OpenStudy (anonymous):

I don't think so

OpenStudy (anonymous):

because when n=1, I get 2!/x, and not 1/x

OpenStudy (xapproachesinfinity):

the summation is n!/x^n for n=0 0!/1=1 for n=1 you have 1!/x=1/x for n=2 you have 2!/x^2=2/x^2

OpenStudy (anonymous):

oh true lol, I did n=-1, and added 1 to the power of x on the bottom, and 1 to the (n+1), so just based on that I should have caught myself....

OpenStudy (xapproachesinfinity):

yeah i see your mistake!

OpenStudy (anonymous):

\[\large \frac{e^x}{x}\left(\begin{matrix} \\ \end{matrix} \sum_{w=0}^{\infty} \frac{w!}{x^{w}} \right)\]

OpenStudy (anonymous):

there it is:)

OpenStudy (xapproachesinfinity):

:) great job pal. when i first saw you integration by part i thought it would involve combinations N choose K but that wasn't the case lol i saw something similar before that turned like that

OpenStudy (anonymous):

yeah, it actually looks neat with the sigma. I feel myself really wise with something like\[\frac{e^x}{x}\left(\begin{matrix} \\ \end{matrix} \sum_{j=0}^{\infty} \frac{(j)!}{x^{j}} \right)\]

OpenStudy (anonymous):

thank you !

OpenStudy (xapproachesinfinity):

yeah it those stuff feel amazing when you arrive at such result welcome pal

OpenStudy (anonymous):

:)